# Yuri Balashov

### v1

@@ 8.3 Opacity and Interpretability: Translation as a Black Box
Despite these considerations, one must acknowledge the overwhelming opacity of how LLMs actually translate. Even if we posit two mechanisms, they are not cleanly separable modules inside the network; they are interwoven in gazillions of parameters. Even if we agree that LLMs’ translation abilities originate from an iterative interplay between two process, Local and Global, the task of saying “how” the model translates a given sentence remains extremely difficult, a familiar problem in explainable AI. For translators and linguists used to working with rule-based or even phrase-based systems, this opacity may be unsettling. It harkens back to the late 20th-century debates on whether translation equivalence can be formally defined or whether it’s a fundamentally intuitive, context-bound judgment (Quine’s indeterminacy of translation comes to mind). With LLMs, the translation process is distributed across layers of matrix multiplications. The duality hypothesis gives us a conceptual handle, but it remains a theoretical abstraction until we can open the black box. Some efforts in interpretability, as we have seen (Section 5.4), are trying to do exactly this—identifying circuits or neurons responsible for aspects of translation. But even if we map some components, the holistic behavior might evade full detailed understanding. Does this matter? In practical terms, maybe not: if the model works, users and developers might not care whether it leverages those aspects of its abilities that were learned via route A rather than B internally. Yet, for trustworthiness and further improvement, it does matter. For instance, if we know a model is relying too much on local memory, we might worry about data bias or inappropriate translations being regurgitated (e.g., if it memorized a biased translation of a term from a specific source). Conversely, if it relies on global semantic alignment, we might be concerned about hallucinations or meaning shifts, especially for critical texts (like medical equipment manuals where a “creative” translation could be dangerous). Thus, understanding the balance between local and global learning is important for governance of MT systems.
@@
It also connects to a deeper point: LLM translation challenges the notion of a translator’s intentionality. Human translators can explain why they chose a phrase (even if post facto, as in the older TAP studies). What about an LLM? We might program it to output a rationale, but that rationale is just another generated text, not an introspective report. This gap has methodological weight as it forces us to ask: can we ascribe concepts like “strategy” or “preference” to an algorithm? The dualism hypothesis suggests that we use such terms by analogy (we can say the model “leans on” one source of knowledge or another), but ultimately these are metaphors. The reality might be that the model has one huge strategy: next-token prediction, and everything else (including translation behaviors) are subsumed under that. We must be careful, then, in attributing human-like decision-making to LLMs. Instead, what we are really observing is a kind of emergent effect that may only resemble having two strategies. This recognition should remind us that any reconceptualization of translation via LLMs has to grapple with the pervasive inscrutability of deep learning. Our theories (duality included) are attempts to impose understandable structure on something intrinsically complex.
@@ 8.4 Continuities with Translation Studies and Distributional Semantics
Retrospectively, it may still be enlightening to place these developments in the context of major turns in translation studies and the rise of distributional semantics in NLP. Translation studies over decades have seen the linguistic turn (focus on equivalence and structure), the cultural turn (focus on context, power, ideology, per Venuti (2013) and other scholars), and more recently a technological turn (examining how tools and automation shape translation, cf. Rothwell et al. 2023; Moorkens et al. 2025). The dual-origin story intersects with all three in surprising ways. The local learning aspect of LLMs is akin to a linguistic approach: it deals with correspondences between words and phrases, almost like a probabilistic dictionary. The global learning aspect echoes the cultural/functional approach: it’s about conveying meaning and intent across languages, not tied to form. Meanwhile, the very existence of LLM translators is itself a result of the technological turn which encourages scholars to revisit longstanding assumptions. For instance, translation quality used to be discussed in terms of equivalence or acceptability, but now we see LLM outputs that are fluent yet may conceal odd source-target mapping behavior (like untranslated bits that “slip through” because the model’s training data had code-switching). This challenges evaluators to differentiate between surface fluency and genuine fidelity. It also rekindles an old debate: Is there a universal translation method? Early MT researchers in the 1950s (following Weaver’s memorandum) searched for a universal code or interlingua; statistical MT later said “no, it’s all about data: the more data, the better”; now LLMs suggest “actually, both approaches co-exist.” This is oddly reminiscent of the compromise proposed by some 20th-century translation theorists who argued for a mixed approach: that translators should sometimes be literal, sometimes free, as needed. It’s as if the pendulum swung and finally settled in the middle within the AI itself.
@@
In conclusion of this section, the age of deep learning invites us—translators, researchers, scholars—to reconceive translation beyond the human-centric, deterministicframeworks. We must account for multiple co-existing methods (even within one model), accept emergence and opacity as part of the game, and borrow insights from both the humanities and computational sciences. This pluralistic, interdisciplinary understanding is not just academic; it will shape how we build and trust future translation systems, and how we train human translators alongside them (e.g., teaching when to trust the “memory” of an LLM vs. when to rely on its “reasoning,” analogous to using a bilingual dictionary vs. paraphrasing an idea).
@@ 9 Conclusion
Another open problem lies in extending high-quality translation to low-resource languages and domains. If our dual-origin hypothesis is correct, it suggests two routes to improvement: feed the model more parallel data for those low-resource cases (to boost local learning) and/or find ways to enhance its semantic alignment (perhaps via continued multilingual training or smarter tokenization that links languages). As we have seen (Sections 4 and 5; see also Xu et al. 2025), there is ongoing work in both directions. An interesting research avenue is to see if synthetic parallel data generated by an LLM (using its global knowledge) can then be used to iteratively improve the same model, essentially bootstrapping global into local knowledge. This could help address the data scarcity issue in a virtuous cycle, but caution is needed to avoid reinforcing errors (a risk if the model’s initial translations are imperfect).

### v3

@@ 8.3 Opacity and Interpretability: Translation as a Black Box
Despite these considerations, one must acknowledge the overwhelming opacity of how LLMs actually translate. Even if we posit two mechanisms, they are not cleanly separable modules inside the network; they are interwoven in gazillions of parameters. Even if we agree that LLMs’ translation abilities originate from an iterative interplay between two process, Local and Global, the task of saying “how” the model translates a given sentence remains extremely difficult, a familiar problem in explainable AI. For translators and linguists used to working with rule-based or even phrase-based systems, this opacity may be unsettling. It harkens back to the late 20th-century debates on whether translation equivalence can be formally defined or whether it is a fundamentally intuitive, context-bound judgment. Quine’s indeterminacy of translation again comes to mind. In a way, the computational opacity of LLMs embodies a version of his underdetermination in mechanistic form. Many distinct internal parameterizations can implement the same input-output behavior (functional equivalence up to reparameterization). With LLMs, the translation process is distributed across layers of matrix multiplications. The duality hypothesis gives us a conceptual handle, but it remains a theoretical abstraction until we can open the black box. Some efforts in interpretability, as we have seen (Section 5.4), are trying to do exactly this—identifying circuits or neurons responsible for aspects of translation. But even if we map some components, the holistic behavior might evade full detailed understanding. Does this matter? In practical terms, maybe not: if the model works, users and developers might not care whether it leverages those aspects of its abilities that were learned via route A rather than B internally. Yet, for trustworthiness and further improvement, it does matter. For instance, if we know a model is relying too much on local memory, we might worry about data bias or inappropriate translations being regurgitated (e.g., if it memorized a biased translation of a term from a specific source). Conversely, if it relies on global semantic alignment, we might be concerned about hallucinations or meaning shifts, especially for critical texts (like medical equipment manuals where a “creative” translation could be dangerous). Thus, understanding the balance between local and global learning is important for governance of MT systems.
@@
It also connects to a deeper point: LLM translation challenges the notion of a translator’s intentionality. Human translators can explain why they chose a phrase (even if post facto, as in the older think-aloud protocol studies; see e.g. bernardini_think-aloud_2001). What about an LLM? We might program it to output a rationale, but that rationale is just another generated text, not an introspective report. This gap has methodological weight as it forces us to ask: can we ascribe concepts like “strategy” or “preference” to an algorithm? The dualism hypothesis suggests that we use such terms by analogy (we can say the model “leans on” one source of knowledge or another), but ultimately these are metaphors. The reality might be that the model has one huge strategy: next-token prediction, and everything else (including translation behaviors) are subsumed under that. We must be careful, then, in attributing human-like decision-making to LLMs. Instead, what we are really observing is a kind of emergent effect that may only resemble having two strategies. This recognition should remind us that any reconceptualization of translation via LLMs has to grapple with the pervasive inscrutability of deep learning. Our theories (duality included) are attempts to impose understandable structure on something intrinsically complex.
@@ 8.4 Continuities with Translation Studies and Distributional Semantics
Retrospectively, it may still be enlightening to place these developments in the context of major turns in translation studies and the rise of distributional semantics in NLP. Translation studies over decades have seen the linguistic turn (focus on equivalence and structure), the cultural turn (focus on context, power, ideology, per venuti_translation_2013 and other scholars), and more recently a technological turn (examining how tools and automation shape translation, cf. rothwell_translation_2023; moorkens_automating_2025). The dual-origin story intersects with all three in surprising ways. The local learning aspect of LLMs is akin to a linguistic approach: it deals with correspondences between words and phrases, almost like a probabilistic dictionary. The global learning aspect echoes the cultural/functional approach: it’s about conveying meaning and intent across languages, not tied to form. Meanwhile, the very existence of LLM translators is itself a result of the technological turn which encourages scholars to revisit longstanding assumptions. For instance, translation quality used to be discussed in terms of equivalence or acceptability, but now we see LLM outputs that are fluent yet may conceal odd source-target mapping behavior (like untranslated bits that “slip through” because the model’s training data had code-switching). This challenges evaluators to differentiate between surface fluency and genuine fidelity. It also rekindles an old debate: Is there a universal translation method? Early MT researchers in the 1950s (following Weaver’s memorandum) searched for a universal code or interlingua; statistical MT later said “no, it’s all about data: the more data, the better”; now LLMs suggest “actually, both approaches co-exist.” This is oddly reminiscent of the compromise proposed by some 20th-century translation theorists who argued for a mixed approach: that translators should sometimes be literal, sometimes free, as needed. It’s as if the pendulum swung and finally settled in the middle within the AI itself.
@@
In conclusion of this section, the age of deep learning invites us—translators, researchers, scholars—to reconceive translation beyond the traditional human-centric frameworks. We must account for multiple co-existing methods (even within one model), accept emergence and opacity as part of the game, and borrow insights from both the humanities and computational sciences. This pluralistic, interdisciplinary understanding is not just academic; it will shape how we build and trust future translation systems, and how we train human translators alongside them (e.g., teaching when to trust the “memory” of an LLM vs. when to rely on its “reasoning,” analogous to using a bilingual dictionary vs. paraphrasing an idea).
@@ 9 Conclusion
Another open problem lies in extending high-quality translation to low-resource languages and domains. If our dual-origin hypothesis is correct, it suggests two routes to improvement: feed the model more parallel data for those low-resource cases (to boost local learning) and/or find ways to enhance its semantic alignment (perhaps via continued multilingual training or smarter tokenization that links languages). As we have seen (Sections 4 and 5; see also xu_survey_2025), there is ongoing work in both directions. An interesting research avenue is to see if synthetic parallel data generated by an LLM (using its global knowledge) can then be used to iteratively improve the same model, essentially bootstrapping global into local knowledge. This could help address the data scarcity issue in a virtuous cycle, but caution is needed to avoid reinforcing errors (a risk if the model’s initial translations are imperfect).
@@
Acknowledgments: The author thanks Chris Wendler, Ryan Nefdt, Michael Carl, members of the Philosophy and Cognitive Science of Deep Learning group, the audience at the AMTA 2025 Virtual: Emerging AI Breakthroughs and Challenges in Translation Automation, the students in the graduate seminar on Language Translation Technologies: Past, Present, and Future taught at the University of Georgia in Fall 2025, and the reviewers for Information for the very helpful feedback on earlier drafts.
Funding: This research was funded by the National Science Foundation Grant No. SES-2336713.
Institutional Review Board Statement: Not applicable.
Informed Consent Statement: Not applicable.
Data Availability Statement: There is no data associated with this research.
Conflicts of Interest: The author declares no conflict of interest.